3 mols of A and 1 mol of B are mixed in a 1L container. the reacion taking place is A=B=2C. if 1.5 mol of C is formed at eqbm.Kc of reaction is 

Grade:12

1 Answers

Khimraj
7 years ago
Intial mole of A B and C are 3,1 and 0. In between reaction are 3-x, 1-x and 2x. And as given 2x = 3/2. So x is 3/4 Then at equilibrium more of A, B and C are 9/4, 1/4 and 3/2 so equilibrium constant is (3/2)²/(9/4)(1/4) = 4. Hope it clears. If u have doubts then please clearify. And if u like my answer then please approve answer